# The Internal Model Principle of Time-Varying Optimization

### v1

@@ III The parameter-feedback problem
Let Assumptions 1, 2, and 3 hold, and let $X_{\circ}\times\Theta_{\circ}$ be some neighborhood of $(x_{\circ}^{\star},0).$ Further, assume that:
@@
Then, there exits a $C^{0}$ mapping $H_{c}:\Theta\rightarrow{}^{n}$ that zeros the gradient locally. $\square$
@@
Then, the result follows from the implicit function theorem (see, e.g., [33, Thm. 1]) applied to the first-order optimality conditions $0=F(x,\theta)$ at the point $(x,\theta)=(x^{\star}_{\circ},0)$. ∎
Intuitively, when the loss function is twice continuously differentiable in the decision variable with a positive definite Hessian, every critical point is a local isolated minimum and existence of $H_{c}(\theta(t))$ is guaranteed by the implicit function theorem. Notice that, differently from the characterization in Theorem 1, the conditions offered by Proposition 2 are sufficient but not necessary.
Although the conditions are immediate to verify, the convergence claims of Theorem 1 and Proposition 2 are of local nature, namely, $y(t)\rightarrow 0$ is ensured provided that $\theta(0)$ is sufficiently close to the origin. The following result provides a sufficient condition for global convergence.
@@
Let Assumptions 1, 2, and 3 hold. Further, assume:
@@
Then, there exits a $C^{0}$ mapping $H_{c}:\Theta\rightarrow{}^{n}$ that zeros the gradient globally. $\square$
@@
and the mappings $\frac{\partial F}{\partial x}$ and $\frac{\partial^{2}F}{\partial x^{2}}$ are $C^{0}$ in ×${}^{n}\times\Theta$. Hence, assumptions (B1)–(B5) of a global version of the implicit function theorem [33, Thm. 2] (see also [34, Thm. 6]) are satisfied for the first-order optimality conditions $0=F(x,\theta)$, and the claim follows. ∎
Theorem 3 shows that, under additional continuity assumptions on the loss function (cf. conditions (i), (ii), and (iv)) and when the Hessian of the loss is positive definite everywhere, convergence of (III) is ensured globally. More formally, under the assumptions of Theorem 3, the parameter-feedback algorithm $x(t)=H_{c}(\theta(t))$ ensures that $y(t)\rightarrow 0$ as $t\rightarrow\infty$ for all initial conditions $\theta(0)\in\Theta$. This result is particularly relevant, as it ensures the existence of parameter-feedback algorithms for the (very important in practice) class of strictly convex loss functions (studied in, e.g., [11, 12]).
We conclude this section by illustrating the design procedure for parameter-feedback algorithms on a quadratic problem.
Example 3.
Consider an instance of (1) with quadratic cost and time-variability that depends linearly on $\theta(t)$ (which has been investigated in [12]):
with matrices $R\in\mathbb{S}^{n}$ and $Q\in{}^{n\times p}$. In this case, the signal we wish to regulate to zero is: $y(t)=\nabla_{x}f(x(t),\theta(t))=Rx(t)+Q\theta(t).$ For arbitrary $\theta,$ this problem admits a critical point if and only if $\operatorname{Im}Q\subseteq\operatorname{Im}R$, in which case $x_{\circ}^{\star}$ is unique. Applying Theorem 1 amounts to finding a linear transformation $H_{c}\in{}^{n\times p}$ such that $0=(RH_{c}+Q)\theta$ for all $\theta$ in a neighborhood of the origin. Assuming $\operatorname{Im}Q\subseteq\operatorname{Im}R$, we can choose $H_{c}=-R^{\dagger}Q$, where $R^{\dagger}$ is the pseudo-inverse of $R$. Note that, by substituting into (III), we have
$y(t)=RH_{c}\theta(t)+Q\theta(t)=0,\quad\forall t\in\mathbb{R}_{\geq 0}.$
Namely, the gradient is identically zero at all times. We also note that this is a particular feature of the parameter feedback problem, which can be achieved because $\theta(t)$ is assumed to be known when implementing (8). As we will see shortly below (cf. Section IV), $y(t)=0$ can be achieved only as $t\rightarrow\infty$ for the gradient-feedback problem.
We conclude by noting that $y(t)=0$ for all $t\in\mathbb{R}_{\geq 0}$ holds for all initial conditions $\theta(0)$ because the mapping $H_{c}(\theta)$ derived here zeros the gradient globally (and not only in some neighborhood of the origin). See also Theorem 3. $\square$
IV The dynamic gradient-feedback problem
In this section, we will build on the framework derived for the parameter-feedback problem to address the main objectives of this work: Problem 1.
IV-A Fundamental results
We begin with the following instrumental characterization.
@@ IV-A Fundamental results
Suppose Assumptions 1, 2, and 3 hold, and assume that $F_{c}(z,y)$ and $G_{c}(z)$ are such that the equilibrium $z=z^{\star}_{\circ}$ of
$\dot{z}(t)$ $=F_{c}(z(t),\nabla_{x}f(G_{c}(z(t)),\theta(t))),$
is exponentially stable. Then, the gradient-feedback optimization algorithm (II-A) asymptotically tracks a critical trajectory of (1) if and only if there exists a $C^{2}$ mapping $z=\sigma(\theta)$ with $\sigma(0)=z_{\circ}^{\star}$, defined in a neighborhood $\Theta_{\circ}\subset\Theta$ of the origin, that satisfies the following for all $\theta\in\Theta_{\circ}$:
$\frac{\partial\sigma(\theta)}{\partial\theta}s(\theta)$ $=F_{c}(\sigma(\theta),0),$ (13a)
00 0 $=\nabla_{x}f(G_{c}(\sigma(\theta)),\theta).$ (13b)
@@
(Only if). We first prove that $\lim_{t\rightarrow\infty}y(t)=0$ implies (13). The coupled dynamics (II-A) have the form:
$\dot{z}$ $=(A_{c}+B_{c}RM)z+B_{c}Q\theta+\chi(x,\theta),$
$\dot{\theta}$ $=S\theta+\psi(\theta).$ (14)
for some mappings $\chi(x,\theta)$ and $\psi(\theta)$ that vanish at the origin along with their first-order derivatives, and where $Q$ and $S$ are defined in (7) and
@@
By assumption, the eigenvalues of the matrix $A_{c}+B_{c}RM$ are in $\mathbb{C}^{-}$ and those of $S$ on the imaginary axis. By Theorem 8, the system (IV-A) has a center manifold at (0,0): the graph of a mapping $z=\sigma(\theta),$ with $\sigma(\theta)$ satisfying (see (31))
$\frac{\partial\sigma(\theta)}{\partial\theta}s(\theta)=F_{c}(\sigma(\theta),\nabla_{x}f(G_{c}(\sigma(\theta)),\theta)).$
By Assumption 3, no trajectory on this manifold converges to zero, so $\lim_{t\rightarrow\infty}y(t)=0$ can hold only if (13b) holds, in which case the above equation reduces to (13a).
(If). We now prove that (13) implies ${\lim_{t\rightarrow\infty}y(t)=0}$. By Theorem 9, the center manifold $z=\sigma(\theta)$ is locally attractive; namely, $z(t)\rightarrow\sigma(\theta(t))$ as $t\rightarrow\infty.$ Then, the fulfillment of (13b) guarantees that $y(t)\rightarrow 0.$
The two conditions in (13) fully characterize the class of optimization algorithms that achieve asymptotic tracking of a critical trajectory. In words, (4) tracks a critical trajectory if and only if, for some mapping $\sigma,$ the composite function $G_{c}\circ\sigma$ zeros the gradient locally (see (13b)), and the controller $F_{c}(z,y)$ is algebraically related to the exosystem $s(\theta)$ as given by (13a). Notice that, by Theorem 1, the former condition implies that
@@
is a parameter-feedback optimization algorithm for (1).
@@
We interpret condition (13a) as the internal model principle of time-varying optimization, as it expresses the requirement that any optimization algorithm that achieves asymptotic tracking must include an internal model of the exosystem. Note that the use of a copy of the temporal variability of the optimization problem is explicit in the prediction-correction algorithm [10] (precisely, through the term $\nabla_{xt}f_{0}(x(t),t)$ – see Remark 1 for notation). $\square$
It is important to note that, by Theorem 4, the exosystem state $\theta$ and that of the optimization $z$ must be related, everywhere in $\Theta_{\circ},$ by the relationship:
@@
Intuitively, (16) is interpreted as the existence of a change of coordinates between the state of the exosystem and that of the optimization (we discuss the invertibility properties of $\sigma(\theta)$ shortly below; see Section IV-B and Remark 4).
Remark 4.
An important special case is obtained when $\sigma$ is the identity operator on $\Theta$; in this case, (13) simplifies to:
$s(\theta)$ $=F_{c}(\theta,0),$
which states that the controller vector field $F_{c}(z,y)$ must coincide with that of the exosystem $s(\theta)$ in $\Theta_{\circ}.$ In this case, (16) gives $z(t)=\theta(t);$ namely, the controller state $z(t)$ and that of the exosystem $\theta(t)$ coincide in $\Theta_{\circ}.$square
While Theorem 4 provides a full characterization of all gradient-feedback algorithms that achieve tracking, it remains to address under what conditions on the loss $f(x,\theta)$ such an algorithm is guaranteed to exist. This question is addressed by the following result.
@@
Suppose Assumptions 1, 2, 3, and 4 hold. There exists a gradient-feedback optimization algorithm that solves Problem 1 if and only if there exists a mapping $H_{c}:\Theta\to{}^{n}$ that zeros the gradient locally. $\square$
@@
(If) We will prove this claim by constructing a gradient-feedback algorithm that achieves $y(t)\rightarrow 0$ as $t\to\infty$.
First, notice that by Assumption 4, there exists a matrix $L$ such that $S-LQ$ has eigenvalues in $\mathbb{C}^{-}.$ Consider the algorithm (4) with $n_{c}=p$ and [^3] Note that the second argument of the gradient is evaluated at the algorithm state $z$ instead of the parameter vector $\theta$ as the latter is unknown.
$F_{c}(z,y)$ $=s(z)+L(y-\nabla_{x}f(H_{c}(z),z)),$
$G_{c}(z)$ $=H_{c}(z),$
where $H_{c}(z)$ is as in (10). The Jacobian of $\dot{z}(t)=F_{c}(z(t),y(t))$ with respect to $z$ is given by $S-LQ$ (notice that two terms of the form $LRM$ with opposite sign cancel out in forming the Jacobian: one from the gradient term $\nabla_{x}f(H_{c}(z),z)$ and the other one from $y$). The claim thus follows by application of Theorem 4 with $\sigma$ the identity operator on $\Theta$. ∎
Interestingly, the conditions for existence of a gradient-feedback algorithm and those for existence of a parameter-feedback algorithm are identical. This is not surprising, since a parameter-feedback algorithm is assumed to have access to $\theta(t)$ while a gradient-feedback algorithm needs to measure $\theta(t)$ indirectly through $y(t).$ More precisely, as stated by (16), the dynamic state of the controller $z(t)$ acts as alternative representation (i.e., in different coordinates) of the exosystem state $\theta(t),$ while the exploration signal $x(t)$ acts as a parameter-feedback algorithm (see (15)).
The proof of Theorem 5 is constructive, as it provides a design procedure to construct $F_{c}(z,y)$ and $G_{c}(z)$ that constitute a gradient-feedback optimization algorithm. Such a procedure is presented in Algorithm 1, where a Luenberger observer is used to estimate the exosystem state $\theta(t)$ (see line $4$), and a parameter feedback algorithm is then applied to the estimated exosystem state to regulate the gradient to zero (precisely, $G_{c}(z)$ is designed following the approach of Theorem 1 – see line $3$).
Data: $s(\theta),$ $\nabla_{x}f(x,\theta),$ $H_{c}(\theta)$ satisfying (13), Jacobian matrices $Q$ and $S$ in (7)
1 $n_{c}\leftarrow n$;

### v2

@@ 3.2 Sufficient conditions for the existence of a parameter-feedback algorithm
Let Assumptions 12 hold, and let $X_{\circ}\times\Theta_{\circ}$ be some neighborhood of $(x_{\circ}^{\star},0).$ Further, assume that:
@@
Then, there exists a $C^{0}$ mapping $H_{c}:\Theta\rightarrow{}^{n}$ that zeros the gradient on some neighborhood of the origin of $\Theta$. $\square$
@@
Then, the result follows from the implicit function theorem (see, e.g., [50, Thm. 1]) applied to the first-order optimality conditions $0=F(x,\theta)$ at the point $(x,\theta)=(x^{\star}_{\circ},0)$. ∎
In other words, Proposition 2 states that when the loss function is sufficiently smooth and the Hessian is locally positive definite, the existence of $H_{c}(\theta(t))$ is guaranteed by the implicit function theorem. Notice that these conditions are sufficient but not necessary.
Remark 9 (Computing mappings that zero the gradient).
Under the assumptions of Proposition 2, the implicit function theorem gives that the linear function
[…]
is a first-order approximation of a mapping that zeros the gradient on a neighborhood of the origin. $\square$
Although the conditions in Proposition 2 are immediate to verify, the convergence claims of Theorem 1 and Proposition 2 are of local nature, namely, $y(t)\rightarrow 0$ is ensured provided that $\theta(0)$ is sufficiently close to the origin. The following result provides a sufficient condition for global convergence.
@@
Let Assumptions 12 hold. Further, assume:
@@
Then, there exits a $C^{0}$ mapping $H_{c}:\Theta\rightarrow{}^{n}$ that zeros the gradient everywhere in $\Theta.$square
@@
and the mappings $\frac{\partial F}{\partial x}$ and $\frac{\partial^{2}F}{\partial x^{2}}$ are $C^{0}$ in ×${}^{n}\times\Theta$. Hence, assumptions (B1)–(B5) of a global version of the implicit function theorem [50, Thm. 2] (see also [51, Thm. 6]) are satisfied for the first-order optimality conditions $0=F(x,\theta)$, and the claim follows. ∎
Theorem 3 shows that, under additional continuity assumptions on the loss function and when the Hessian of the loss is positive definite everywhere, there exists a mapping $H_{c}:\Theta\rightarrow{}^{n}$ that zeros the gradient everywhere in $\Theta.$
By combination of Theorem 3 and Theorem 1, it follows that under the assumptions of Theorem 3, the parameter-feedback algorithm $x(t)=H_{c}(\theta(t))$ ensures that $y(t)\rightarrow 0$ as $t\rightarrow\infty$ for all initial conditions $\theta(0)\in\Theta$. This result ensures the existence of parameter-feedback algorithms for the particular class of strictly convex loss functions [11, 12].
4 The dynamic gradient-feedback problem
In the previous section, we analyzed algorithms assuming knowledge of the function $\nabla_{x}f(x,\theta)$ and of the signal $\theta(t),t\in\mathbb{R}_{\geq 0}$ (cf. assumptions (E1) and (O2)). In this section, we relax these assumptions and require only that $y(t)$ is available through oracle evaluations of (9). For this reason, we will shift our attention back to the general class of dynamic gradient-feedback algorithms (10).
Because $\theta(t)$ is unmeasurable by the algorithm, the temporal variability of the cost can only be evaluated through measurements of $y(t);$ for this reason, we impose the following.
Assumption 3 (Detectability of the exosystem).
Let
The pair $(Q,S)$ is detectable. $\square$
The detectability property of Assumption 3 is equivalent to the existence of a local exponential observer for all (non-asymptotically stable) modes of $\theta(t)$ based on measurements $y(t)$ [52, Cor. 3.4]. Undetectability corresponds to a redundant description of the exogenous signal: if some modes of $\theta(t)$ are undetectable, these modes do not influence the gradient (and, as a consequence, the critical points of (1)), and thus they can be removed from the optimization problem (1) without altering its critical points. In other words, Assumption 3 ensures that (2) does not contain redundancies.
4.1 Existence and characterization of dynamic gradient feedback algorithms
We begin by considering the class of algorithms (10), restricted to ensure local exponential stability, and we characterize a set of necessary and sufficient conditions for this class to achieve exact asymptotic tracking.
@@ 4.1 Existence and characterization of dynamic gradient feedback algorithms
Suppose Assumptions 13 hold, and assume that $F_{c}(z,y)$ and $G_{c}(z)$ are such that the equilibrium $z=z^{\star}_{\circ}$ of
$\dot{z}(t)$ $=F_{c}(z(t),\nabla_{x}f(G_{c}(z(t)),0)),$ (22)
is locally exponentially stable. The optimization algorithm (12) exactly asymptotically tracks a critical trajectory of (1) if and only if there exists a neighborhood $\Theta_{\circ}\subset\Theta$ of the origin and a $C^{2}$ mapping $z=\sigma(\theta)$ with $\sigma(0)=z_{\circ}^{\star}$, such that:
$\left.\frac{\partial\sigma(\theta)}{\partial\theta}\right|_{\theta=\theta_{\omega}}s(\theta_{\omega})$ $=F_{c}(\sigma(\theta_{\omega}),0),$ (23a)
00 0 $=\nabla_{x}f(G_{c}(\sigma(\theta_{\omega})),\theta_{\omega}),$ (23b)
at all limit points $\theta_{\omega}\in\Omega(\Theta_{\circ})$.
@@
(Only if). We first prove that $\lim_{t\rightarrow\infty}y(t)=0$ implies (23). The coupled dynamics (12) have the form:
$\dot{z}$ $=(A_{c}+B_{c}RM)z+B_{c}Q\theta+\chi(z,\theta),$
$\dot{\theta}$ $=S\theta+\psi(\theta),$ (24)
for some mappings $\chi(x,\theta)$ and $\psi(\theta)$ that vanish at the origin along with their first-order derivatives, with $Q$ and $S$ are defined in (21), and
@@
By assumption, the eigenvalues of the matrix $A_{c}+B_{c}RM$ are in $\mathbb{C}^{-}$. By Theorem 9, the system (4.1) has a center manifold at $(z^{\star}_{\circ},0)$: the graph of a mapping $z=\sigma(\theta),$ with $\sigma(\theta)$ satisfying (see (41))
$\sigma(s(\theta))=F_{c}(\sigma(\theta),\nabla_{x}f(G_{c}(\sigma(\theta)),\theta)).$
Next, similarly to the parameter-feedback case (Theorem 1), by Assumption 2, $\theta(t_{i})$ converges to some limit point $\theta_{\omega}\in\Omega\left(\theta(0)\right)$ for some subsequence $\{t_{i}\}_{i\in\mathbb{N}_{\geq 0}}$. By continuity of the gradient (Assumption 1) and that of $G_{c}$,
$\lim_{i\to\infty}y({k_{i}})=\lim_{i\to\infty}\nabla_{x}f(G_{c}(\theta_{t_{i}}),\theta_{t_{i}})=\nabla_{x}f(G_{c}(\theta_{\omega}),\theta_{\omega}).$ (25)
Because, $y(t)\to 0$, the left-hand side of (25) is equal to zero, and thus $\nabla_{x}f(G_{c}(\theta_{\omega}),\theta_{\omega})=0.$ Since this holds for any limit point $\theta_{\omega}\in\Omega(\theta_{0})$, the claim follows.
(If). We now prove that (23) implies ${\lim_{t\rightarrow\infty}y(t)=0}$. It follows from (23) and Theorem 10 that the graph of the mapping $z=\sigma(\theta)$ (i.e., $(\sigma(\theta),\theta)$) is a center manifold for the coupled dynamics (12). Moreover, such manifold is locally attractive; namely, $z(t)\rightarrow\sigma(\theta(t))$ as $t\rightarrow\infty.$ Then, the fulfillment of (23b) guarantees that the right-hand side of (25) $y(t)\rightarrow 0.$ The conclusion then follows by iterating the (If) part of the proof of Theorem 1. ∎
The two conditions in (23) fully characterize the class of optimization algorithms that achieve exact asymptotic tracking: (10) tracks a critical trajectory if and only if, for some mapping $\sigma,$ the composite function $G_{c}\circ\sigma$ zeros the gradient locally (see (23b)), and the controller $F_{c}(z,y)$ is algebraically related to the exosystem $s(\theta)$ as given by (23a). Notice that, by Theorem 1, the former condition implies that
@@
is a parameter-feedback optimization algorithm for (1). Finally, we note that the existence of an exponentially stable system as in (22) is always guaranteed under Assumption 3 [52, Sec. 3].
The following interpretation follows from our findings.
@@
We refer to condition (23a) as the internal model principle of time-varying optimization, as it encapsulates the requirement that an optimization algorithm must incorporate an internal model of the exosystem to achieve exact asymptotic tracking. Note that the use of a copy of the temporal variability of the optimization problem is explicit in the prediction-correction algorithm (see (7)). $\square$
Theorem 4 allows us to re-interpret the basic gradient-flow algorithm as follows.
Remark 11 (Internal model-based interpretation of basic gradient flow methods).
Recall that the gradient-flow algorithm (4) can be viewed as an instance of (10a) with $F_{c}(z,y)$ and $G_{c}(z)$ as in Remark 7. By direct substitution into (23), it is immediate to see that this algorithm satisfies (23a) with $s(\theta)=0$ and $\sigma(\theta)$ arbitrary. Since $s(\theta)=0$ describes the internal model of a constant signal, it follows from Theorem 4 that these algorithms achieve exact asymptotic tracking if and only if $\theta(t)$ is a constant signal. This observation is in line with the inexact convergence properties of these algorithms typically given for these algorithms [48, 26, 14$\square$
Since by, Theorem 4, the problem of exact asymptotic tracking can be reduced to that of finding an optimization algorithm that is algebraically equivalent to the exosystem and a mapping that zeros the gradient, we have the following.
Answer to Problem 2.2. When an algorithm as in (10) that achieves exact asymptotic tracking exists, the minimal knowledge needed to design such an algorithm is given by assumptions (E2) and (O2). Notice that assumption (O1) is implicitly required to execute the algorithm (10).
By comparison with Answer to Problem 2, it follows that relaxing the knowledge of $\theta(t)$ (i.e., relaxing (E1)) comes at the cost of requiring knowledge of the exosystem (i.e., requiring (E2)).
By Theorem 4, the exosystem state $\theta$ and that of the optimization $z$ must be related, everywhere in $\Theta_{\circ},$ by
@@
Intuitively, (27) is interpreted as the existence of a change of coordinates between the state of the exosystem and that of the optimization; see Section 4.3 for a discussion on the invertibility properties of the function $\sigma(\theta)$.
Remark 12 (Special case with $\sigma$ being the identity operator).
An important special case is obtained when $\sigma$ is the identity operator on $\Theta$; in this case, (23) simplifies to $s(\theta)=F_{c}(\theta,0)$, which states that the controller vector field $F_{c}(z,y)$ must coincide with that of the exosystem $s(\theta)$ in $\Omega(\Theta_{\circ}).$ In this case, (27) gives $z(t)=\theta(t);$ namely, the controller state $z(t)$ and that of the exosystem $\theta(t)$ coincide in $\Omega(\Theta_{\circ}).$square
While Theorem 4 provides a full characterization of all gradient-feedback algorithms that achieve tracking, it remains to address under what conditions such an algorithm is guaranteed to exist. This question is addressed next.
@@
Suppose Assumptions 13 hold. Then, there exists a gradient-feedback algorithm, such that $z=z^{\star}_{\circ}$ is exponentially stable for (22), that solves Problem 1 if and only if there exists a mapping $H_{c}:\Theta\to{}^{n}$ that zeros the gradient on the limit set of (2) with respect to its initializations. $\square$
@@
(If) We will prove this claim by constructing a gradient-feedback algorithm that achieves $y(t)\rightarrow 0$ as $t\to\infty$. First, notice that by Assumption 3, there exists a matrix $L$ such that $S-LQ$ has eigenvalues in $\mathbb{C}^{-}$. Consider the algorithm (10) with $n_{c}=p$ and [^6] Note that the second argument of the gradient is evaluated at the algorithm state $z$ instead of the parameter vector $\theta$ as the latter is unknown.
$F_{c}(z,y)$ $=s(z)+L(y-\nabla_{x}f(H_{c}(z),z)),$ (28a)
$G_{c}(z)$ $=H_{c}(z),$ (28b)
where $H_{c}(z)$ is as in (16). The Jacobian of $\dot{z}(t)=F_{c}(z(t),y(t))$ with respect to $z$ is given by $S-LQ$ (two terms of the form $LRM$ with opposite sign cancel out in forming the Jacobian: one from the gradient term $\nabla_{x}f(H_{c}(z),z)$ and the other one from $y$); since $S-LQ$ has eigenvalues in $\mathbb{C}^{-}$, $z=z^{\star}_{\circ}$ is exponentially stable for (22). The claim thus follows by Theorem 4 with $\sigma$ the identity operator on $\Theta$. ∎
Interestingly, the conditions for the existence of a gradient-feedback algorithm and those for the existence of a parameter-feedback algorithm (cf. Section 3) are identical. This is not surprising, since a parameter-feedback algorithm is assumed to have access to $\theta(t)$ while a gradient-feedback algorithm needs to measure $\theta(t)$ indirectly through $y(t).$ More precisely, as stated by (27), the dynamic state of the controller $z(t)$ acts as an alternative representation (i.e., in different coordinates) of the exosystem state $\theta(t),$ while the exploration signal $x(t)$ acts as a parameter-feedback algorithm (see (26)).
4.2 Design of dynamic gradient feedback algorithms
We now focus on addressing Problem 1. A design procedure to construct $F_{c}(z,y)$ and $G_{c}(z)$ is presented in Algorithm 1, which is derived from the proof of Theorem 5.
The algorithm uses a Luenberger observer to estimate the exosystem state $\theta(t)$ (cf. line $4$), and a parameter feedback algorithm is then applied to the estimated exosystem state to regulate the gradient to zero (precisely, $G_{c}(z)$ is designed following the approach of Theorem 1 – cf. line $3$ of Algorithm 1). Notice that the algorithm uses the particular choice for $\sigma(\theta)$ being the identity operator (see Remark 12). Note that line $2$ of the algorithm guarantees that the equilibrium of (22) is locally exponentially stable, as required by Theorem 4, and that the existence of such $L$ is guaranteed by Assumption 3.
Data: $s(\theta),$ $\nabla_{x}f(x,\theta),$ $G_{c}(\theta)$ satisfying (23b), Jacobian matrices $Q$ and $S$ in (21)
1 $n_{c}\leftarrow p$;